OpenStudy (anonymous):

integrate x^2/(x^2+1)^1/2

OpenStudy (luigi0210):

What have you tried so far?

ganeshie8 (ganeshie8):

whenever you see a radical, the first thing to try is to substitute it : \[\large u^2 = x^2+1\]

ganeshie8 (ganeshie8):

i bet you have tried that already :) it doesn't seem to work out smoothly hmm

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

what about trig substitution ?

OpenStudy (anonymous):

not tried

ganeshie8 (ganeshie8):

try it, looks it simplifies nicely u = tan

OpenStudy (anonymous):

not getting

ganeshie8 (ganeshie8):

are you on words diet today hmm ;p

OpenStudy (anonymous):

may be

OpenStudy (anonymous):

cn u give me the exact solution

ganeshie8 (ganeshie8):

well i am on diet too >.<

OpenStudy (anonymous):

oh so u r not going to help me

ganeshie8 (ganeshie8):

i can try to help if you atleast attempt something

ganeshie8 (ganeshie8):

even i don't know how to work it yet, so we need to work it together

OpenStudy (anonymous):

i dont know about trig. substitution

ganeshie8 (ganeshie8):

how about hyperbolics ?

ganeshie8 (ganeshie8):

or something like this : multiply numerator and denominator by x^2, pull out a x from the botom radical and substitute \[\large u = 1+\frac{1}{x^2}\]

ganeshie8 (ganeshie8):

they are all the things i would try if that was my homework problem

ganeshie8 (ganeshie8):

notice that derivative of arcsinhx is 1/(x^2+1)^1/2 that might help in getting ideas on how to attempt too

OpenStudy (anonymous):

the answer given is in log

OpenStudy (shinalcantara):

Integration by parts perhaps may help

OpenStudy (anonymous):

shinal what should i take as first function then

OpenStudy (shinalcantara):

okay it took me a while \[\int\limits \frac{ x ^{2} }{ (x + 1)^{\frac{ 1 }{ 2 }} }dx\] can be expressed as \[\int\limits (x)(\frac{ x }{ (x ^{2}+1)^{\frac{ 1 }{ 2 }} })\] have you remembered \[\int\limits udv = uv - \int\limits vdu\] let's do that for the given expression we let \[u = x \] \[du = dx\] \[v = \frac{ x }{ (x ^{2}+1)^{\frac{ 1 }{ 2 }} }dx\] or \[dv = x(x ^{2}+1)^{-\frac{ 1 }{ 2 }}dx\] integrate to have the value for v \[\int\limits \frac{ dv }{ dx } = \frac{ 1 }{ 2 } \int\limits 2x(x ^{2}+1)^{-\frac{ 1 }{ 2 }}dx\] \[v = \frac{ 1 }{ 2 }(\frac{ (x^2+1)^{\frac{ 1 }{ 2 }} }{ \frac{ 1 }{ 2 } })\] \[v = (x^2 + 1)^{\frac{ 1 }{ 2 }}\] let's substitute it to \[\int\limits udv = uv - \int\limits vdu\] \[ = x(x^2+1)^{\frac{ 1 }{ 2 }} - \int\limits (x^2+1)^{\frac{ 1 }{ 2 }}dx\] since \[\int\limits \sqrt{u^2+a^2} du = \ln \left[ \sqrt{u^2+a^2}+u \right] +c\] then we'll apply that and we'll have \[= x(x^2+1)^\frac{ 1 }{ 2 } - \ln \left| \sqrt{x^2+1}+x \right| + c\]

OpenStudy (ikram002p):

u still need help :O

OpenStudy (anonymous):

no ,thanks